(-2/3)+p=1/6
show work

Answer by zeynep(43)

(-2/3)+p=1/6 We add additive inverse of (-2/3) to both sides;
(-2/3)+p+(2/3)=1/6+(2/3)
p=5/6
To solve this kind of equations we get the variable alone on one side by using inverse operations.